A 50kg cat statue falls from a second floor (10m high) window. With what momentum does it impact the ground with?


1
Expert's answer
2021-03-17T08:34:27-0400

The impact velocity

"mgh = \frac{mv^2}{2} \\

v = \sqrt{2gh} \\

= \sqrt{2 \times 9.8 \times 10} = 14 \;m/s"

The momentum:

"p = mv \\

= 50 \times 14 = 700 \;kg \times m/s"